OpenStudy (cicilybailey14):

Hey qt's I need some help. What is the square root of 3 to the square root of 2 power times the square root of 3 to the negative square root of 2 power? A. 100 B. 10 C. 1 D. 10^3

OpenStudy (croit):

Could you please write it in "Equation"form.

OpenStudy (mikey):

what does the equation look like its hard to picture with words

OpenStudy (cicilybailey14):

There is no equation with it,,,

OpenStudy (croit):

It doesn't need to be an equation. Just write it like \[x ^{a}^{b}\]...

OpenStudy (mikey):

so the question is all words and no actual square roots signs or exponents

jhonyy9 (jhonyy9):

\[\sqrt{3}^\sqrt{2}*(\sqrt{3})^{-\sqrt{2}}\]

jhonyy9 (jhonyy9):

do you think it same ?

OpenStudy (triciaal):

you can use the draw below click on the pencil when finished remember to post

OpenStudy (triciaal):

apply the rules for exponents when you have the same base

OpenStudy (croit):

@jhonyy9 I think you are right.

OpenStudy (triciaal):

did anyone else get 1?

OpenStudy (croit):

Of course it is equal to 1, if the given expression is the same as given by @jhonyy9 .